OpenStudy (anonymous):

solve for x: logx(3)^1/2 + logx 3^5 + logx 1/27 = 5/4. Can somebody help me.....

OpenStudy (ash2326):

Is this your question? \[\log_x {3^{\frac 12}}+\log_x {3^5}+\log_x \frac{1}{27}\ \ \ \ \ \ =\frac 5 4\]

OpenStudy (anonymous):

yes

OpenStudy (ash2326):

do you know the addition property of logarithms ? \[\log_x a +\log_x b=??\]

OpenStudy (anonymous):

yes, it's logx ab right..

OpenStudy (ash2326):

good :) so here we have \[\log_x {3^{\frac 12}}+\log_x {3^5}+\log_x \frac{1}{27}\ \ \ \ \ \ =\frac 5 4\] \[\log_x {3^{\frac 12}}+\log_x {3^5}+\log_x {3^{-3}}\ \ \ \ \ \ =\frac 5 4\]

OpenStudy (ash2326):

Could you use the addition property to simplify the left side?

OpenStudy (anonymous):

do I multiply the same base? logx (3^1/2)(3^5)(3^-3)

OpenStudy (ash2326):

Base is same, so we can use the addition property \[\log_x {3^{\frac 12}3^{5}3^{-3}}=\frac 54\]

OpenStudy (anonymous):

so it would be logx 3^5/2

OpenStudy (ash2326):

good, now use the fundamental property of log to find x

OpenStudy (anonymous):

3^5/2=x^5/4 ?

OpenStudy (ash2326):

can you create power of 5/4 on the left side

OpenStudy (anonymous):

how do i do that... :S

OpenStudy (ash2326):

\[\huge 3^{\frac 52}=3^{\frac {5\times 2}{2\times 2}=(3^2)^{\frac 54}}\]

OpenStudy (anonymous):

why are we making the power the same.. i dont understand sorry im abit slow

OpenStudy (ash2326):

\[2^4=x^4\] x=2 we need same powers so that we could find x :)

OpenStudy (anonymous):

Very easy question we make powers same to compare and make our work easy

OpenStudy (ash2326):

\[\huge ({3^2})^{\frac 54}=x^{\frac 54}\] so x= ??

OpenStudy (anonymous):

9?

OpenStudy (anonymous):

correct

OpenStudy (ash2326):

yes, it's correct. Do you understand this?

OpenStudy (anonymous):

yes now i do, but usually I cancel the number not the power so it was kind of confusing..

OpenStudy (anonymous):

THANK YOU :D

OpenStudy (anonymous):

indices dont follow same rules of basic maths

OpenStudy (anonymous):

oh it's allowed to cancel the power too?

OpenStudy (anonymous):

if you have same bases for eg

OpenStudy (anonymous):

\[3^2=3^2\]

OpenStudy (anonymous):

or 3^9=y^9 ... so y=3?

OpenStudy (anonymous):

like that..?

OpenStudy (anonymous):

yes correct now ur clear

OpenStudy (anonymous):

yes thanks you :D

OpenStudy (anonymous):

thank*

OpenStudy (anonymous):

welcome for which exam are you studying

OpenStudy (anonymous):

final exam for Foundation in Economics :)

OpenStudy (anonymous):

I am in high School studying for India's most prestigious exam in India IIT

OpenStudy (anonymous):

wow your last year in high school?

OpenStudy (anonymous):

No second last in India we call it as FYJC just search for IIT in google

OpenStudy (anonymous):

it's showing some kind of higher education thing.

OpenStudy (anonymous):

No its graduation BTech

OpenStudy (anonymous):

ohh okay sounds hard tho. All the best! :D

OpenStudy (anonymous):

Thanks for u also all the Best wishes

OpenStudy (anonymous):

der on fb

OpenStudy (anonymous):

Thank you I need that :)